OpenStudy (anonymous):

The product of a number and 13 is no more than 10

OpenStudy (anonymous):

x<=10/13

OpenStudy (ericanoel912):

\[x \times 13 \le 10\]

OpenStudy (ericanoel912):

there's no division in this problem- just multiplication

OpenStudy (texaschic101):

ericanoel is right.....the product is the result of multiplication it can be : n * 13 <= 10 or 13n <= 10 (and your sign is less then or equal)